OpenStudy (anonymous):

A regular square pyramid has a slant height equal to the length of its base edge. a) Write a formula for the surface area in terms of its base edge, e. b) Find the surface area of the pyramid if e=200ft.

OpenStudy (jdoe0001):

hmm do you know what a square pyramid is?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so... I assume you know how to find the area of a triangle?

OpenStudy (jdoe0001):

 so the lateral surface area will be the SUM of the area of each of those 4 triangular faces

OpenStudy (anonymous):

um...okay....so the slant height is the same number as the base?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

or one can say, the slant height of the pyramid, is the same as the base edge or the altitude of the triangular face, is the same as it's base

OpenStudy (anonymous):

wow. ok. so the answer is: 1000ft?

OpenStudy (anonymous):

what formula should I use?

OpenStudy (jdoe0001):

hmmm wel... you're supposed to write it so just write the sum of all triangle's area then simplify it

OpenStudy (anonymous):

so I write: \[1/2 base * height = 100*100 = 1,000feet\]

OpenStudy (jdoe0001):

1,000? where did that come from?

OpenStudy (jdoe0001):

I read that the slant height is just "e"

OpenStudy (jdoe0001):

if your base and slant height is 100, yes

OpenStudy (anonymous):

so the base isn't 200?

OpenStudy (anonymous):

Could you tell me the answer then? Cuz im confused.

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

Is the equation: 200*200/2 ? @agent0smith